A particle of mass 2 kg moves with initial velocity of v=4i+4j. a constant force of F=-20j N is applied on the particle. initially particle was at (0,0) . the X-coordinate of the particle when it's y coordinate again becomes zero is given by

Grade:9

1 Answers

Mohd As an Ali
5 years ago
Let the particle travel for time t,  then coordinate of position after time t will be(x, 0). 
Now F=ma; -20=2•a j;  a=-10jm/s2
as there is no displacement along y direction hence s=0 we use newtons second law of motion. 
S=ut+0.5at2.
0=4t--0.5•10t2
4t=5t2
t=0.8s 
Now displacement along i is x i and velocity is 4i.
Then 4=x/t    ;     x=3.2
 
Hence coordinate is (3.2, 0).